# Investigating the influence of radio-faint AGN activity on the infrared-radio correlation of massive galaxies

### v1

@@ 1 Introduction
Several studies (Magnelli et al., 2015; Rivera et al., 2017; Delhaize et al., 2017) observed a dependence of the IRRC on the redshift, in the form of with . Other works (Gürkan et al., 2018; Smith et al., 2021; Delvecchio et al., 2021, D21 hereafter) have recently argued that the IRRC evolves primarily with , with more massive galaxies showing a systematically lower , while the dependence on the redshift is secondary and weaker, as indeed at fixed stellar mass (see also, Gürkan et al., 2018; Smith et al., 2021). Among the possible physical drivers of the decline observed in massive star-forming galaxies, the energy loss of cosmic ray electrons (CRe) has been suggested to play a significant role (e.g. Schober et al., 2023; Yoon, 2024). Specifically, Yoon (2024) modelled the CRe energy-loss as a function of the gas density and redshift of the host galaxy. Their findings indicate an anti-correlation between these quantities: specifically, galaxies with higher gas densities (or, equivalently, higher stellar masses, e.g. Kennicutt, 2003), experience less intense energy losses, arguably because they retain more easily their CRes. On the reverse, less massive galaxies are generally most affected by CRe’s energy losees. As a consequence of these synchrotron energy losses, less massive galaxies have lower radio luminous with respect to more massive ones.
@@
where is the observed density flux at the frequency and and are the major and minor axes of the beam (Ulvestad et al., 2004).
@@ 3 Observation and data calibration
where is the number of ’real’ (input) blobs and is the number of the (output) detected blobs. We count the output blobs within a radius of 0.2′′ from the input (real) source, similarly to what we do for real detections (see Sect.3.1). We find that the spuriousness is higher than 50% at S/N 5, while it decreases to 15% at S/N5.5, and exponentially drops below 2% at S/N6. If counting all random blobs over the full image (44′′), the spuriousness at fixed S/N increases proportionally to the searching area, except at S/N6 where it drops to below 2% as found in the central area.
3.1 Final sample of VLBA detections
@@
represents the detections, and the hatched orange histogram is the so-called mirror distribution, interpreted as the intrinsic distribution of SFGs. The bootstrapped median values of the mirror distribution were fitted with a Gaussian (red curve; see Sec. 4 for further details). Left panel: is computed using the VLA luminosity coming from both star formation and AGN activity. Right panel: is the ratio measured after subtracting the AGN radio luminosity detected by the VLBA from the total radio VLA emission.
@@
Then, we also find that the real distribution exhibits a deviation from the Gaussianity (e.g., when compared to the mirror distribution) on the left side, in both panels. However, this is minimal, as indeed
a K-S test could not reject the hypothesis that these two distributions come from the same parent sample (with a p-value of 0.997).
@@ 6.1 Cumulative Distribution of AGN in the faintest regime
To do so, we convert the Euclidean-normalized 1.4 GHz source counts into a dimensionless quantity, re-scaled to the effective sky area of the (,rms)-matched CS (i.e. 9.7 x 10-6 sr). This yields a number of expected AGN in the same area, as a function of 1.4 GHz flux. We make sure that the relation between differential NAGN and S1.4 yields an integrated = 4 in the range 36–58 Jy (i.e. matching our VLBA detections, red stars in Fig. 4). Then we extrapolate the same trend down, and below our nominal VLBA flux density limit. Finally, we compute the cumulative distribution of at Jy, and we found out that 42 sources would be cumulated at a flux density of 5 Jy, which is then the minimum flux density reached by the 42 brightest AGN. We stress again that this is an over-conservative assumption, since it is very unlikely that our non-detections are all clustered just below the flux limit. However, at these low flux densities, the brightness temperature drops to K (see Eq. 2). This suggests that, even if we were able to detect such faint sources with the VLBA, the origin of their emission would be ambiguous between AGN or SF. Consequently, we conclude that pushing VLBA observation to sub-Jy sensitivity would not be an effective strategy for revealing a residual hidden population of radio AGN. For instance, according to the ngVLA exposure time calculator 555https://ngect.nrao.edu/, we note that Band 1 (2.3 GHz observations) would reach brightness temperatures T K (at 5, i.e. 5 Jy/beam) in about 16h on-source. Hence, fainter fluxes, as expected for the bulk of our VLBA non-detections, would be detected with a lower T K, thus carrying an ambiguous origin despite their easy detection.
@@ 6.2 Diffuse AGN emission as a possible caveat
Assuming that this radio-excess traces the full AGN-related flux over the entire galaxy size, the comparison between
total VLBA and AGN-driven VLA flux indicates how much flux is lost on larger scales.
We note (see Figure 11 in Herrera-Ruiz et al., 2017) that at fluxes S1.4 , the VLBA flux is already of the total VLA AGN flux, hence highly compact. By extrapolating this trend towards fainter flux regimes (as for our VLBA targets), we would expect that the VLBA flux is capturing the vast majority (if not all) of the total AGN flux.
@@ Acknowledgements.
The authors thanks the referee for a constructive report that helped us clarify the results and implications of this work. G. P., I. D., M. G., F. U., and I. P. acknowledge funding by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006. C. S. acknowledges financial support from INAF Ricerca Fondamentale 2024 (Ob.Fu. 1.05.24.07.04), from the Italian Ministry of Foreign Affairs and International Cooperation (grant PGRZA23GR03), and by the Italian Ministry of University and Research (grant FIS-2023-01611, CUP C53C25000300001). W.W. acknowledges the grant support through JWST programs. Support for programs JWST-GO-03045 and JWST-GO-03950 were provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127.References
@@
- Algera et al. (2020) Algera, H. S. B., Smail, I., Dudzevičiūtė, U., et al. 2020, The Astrophysical Journal, 903, 138
- Appleton et al. (2004) Appleton, P. N., Fadda, D. T., Marleau, F. R., et al. 2004, The Astrophysical Journal Supplement Series, 154, 147
- Basu et al. (2015) Basu, A., Wadadekar, Y., Beelen, A., et al. 2015, Astrophysical Journal, 803, 51
- Bell (2003) Bell, E. F. 2003, The Astrophysical Journal, 586, 794
- Bourne et al. (2011) Bourne, N., Dunne, L., Ivison, R. J., et al. 2011, Monthly Notices of the Royal Astronomical Society, Volume 410, Issue 2, pp. 1155-1173., 410, 1155
- Casey et al. (2023) Casey, C. M., Kartaltepe, J. S., Drakos, N. E., et al. 2023, The Astrophysical Journal, 954, 31
@@
- Condon (1992) Condon, J. J. 1992, Annual Rev. Astron. Astrophys., Vol. 30, p. 575-611 (1992), 30, 575
@@
- Deane et al. (2024) Deane, R. P., Radcliffe, J. F., Njeri, A., et al. 2024, Monthly Notices of the Royal Astronomical Society, Volume 529, Issue 3, pp.2428-2442, 529, 2428
@@
- Delvecchio et al. (2022) Delvecchio, I., Daddi, E., Sargent, M. T., et al. 2022, Astronomy and Astrophysics, 668, A81
- Delvecchio et al. (2021) Delvecchio, I., Daddi, E., Sargent, M. T., et al. 2021, Astronomy and Astrophysics, 647, A123
- Delvecchio et al. (2017) Delvecchio, I., Smolčić, V., Zamorani, G., et al. 2017, Astronomy and Astrophysics, 602, A3
@@
- Galluzzi et al. (2025) Galluzzi, V., Behiri, M., Giulietti, M., & Lapi, A. 2025, eprint arXiv:2504.00192, arXiv:2504.00192
@@
- Gim et al. (2019) Gim, H. B., Yun, M. S., Owen, F. N., et al. 2019, The Astrophysical Journal, 875, 80
- Gürkan et al. (2018) Gürkan, G., Hardcastle, M. J., Smith, D. J., et al. 2018, Monthly Notices of the Royal Astronomical Society, 475, 3010
- Hales et al. (2012) Hales, C. A., Murphy, T., Curran, J. R., et al. 2012, Monthly Notices of the Royal Astronomical Society, 425, 979
- Harish et al. (2025) Harish, S., Kartaltepe, J. S., Liu, D., et al. 2025, arXiv, arXiv:2506.03306
- Harwit & Pacini (1975) Harwit, M. & Pacini, F. 1975, Astrophysical Journal, vol. 200, Sept. 15, 1975, pt. 2, p. L127-L129., 200, L127
- Helou & Rowan-Robinson (1985) Helou, G. & Rowan-Robinson, M. 1985, The Astrophysical Journal, 298, 7
- Herrera-Ruiz et al. (2017) Herrera-Ruiz, N., Middelberg, E., Deller, A., et al. 2017, Astronomy and Astrophysics, 607, A132
- Herrera-Ruiz et al. (2018) Herrera-Ruiz, N., Middelberg, E., Deller, A., et al. 2018, Astronomy and Astrophysics, 616, A128
- Heywood et al. (2022) Heywood, I., Rammala, I., Camilo, F., et al. 2022, The Astrophysical Journal, 925, 165
@@
- Ivison et al. (2010a) Ivison, R. J., Alexander, D. M., Biggs, A. D., et al. 2010a, Monthly Notices of the Royal Astronomical Society, 402, 245
- Ivison et al. (2010b) Ivison, R. J., Magnelli, B., Ibar, E., et al. 2010b, Astronomy and Astrophysics, 518, L31
- Jarvis et al. (2010) Jarvis, M. J., Smith, D. J. B., Bonfield, D. G., et al. 2010, Monthly Notices of the Royal Astronomical Society, Volume 409, Issue 1, pp. 92-101., 409, 92
@@
- Jin et al. (2018) Jin, S., Daddi, E., Liu, D., et al. 2018, The Astrophysical Journal, 864, 56
- Kartaltepe et al. (2015) Kartaltepe, J. S., Mozena, M., Kocevski, D., et al. 2015, Astrophysical Journal, Supplement Series, 221, 11
- Kashino et al. (2019) Kashino, D., Silverman, J. D., Sanders, D., et al. 2019, The Astrophysical Journal Supplement Series, 241, 10
- Kennicutt (2003) Kennicutt, R. C. 2003, https://doi.org/10.1146/annurev.astro.36.1.189, 36, 189
@@
- Madau & Dickinson (2014) Madau, P. & Dickinson, M. 2014, Annual Review of Astronomy and Astrophysics, 52, 415
- Magnelli et al. (2015) Magnelli, B., Ivison, R. J., Lutz, D., et al. 2015, Astronomy and Astrophysics, 573, A45
@@
- Middelberg et al. (2013) Middelberg, E., Deller, A. T., Norris, R. P., et al. 2013, Astronomy and Astrophysics, 551, A97
@@
- Molnár et al. (2018) Molnár, D. C., Sargent, M. T., Delhaize, J., et al. 2018, Monthly Notices of the Royal Astronomical Society, 475, 827
- Molnár et al. (2021) Molnár, D. C., Sargent, M. T., Leslie, S., et al. 2021, Monthly Notices of the Royal Astronomical Society, 504, 118
@@
- Njeri et al. (2023) Njeri, A., Beswick, R. J., Radcliffe, J. F., et al. 2023, Monthly Notices of the Royal Astronomical Society, 519, 1732
@@
- Ponnada et al. (2025) Ponnada, S. B., Cochrane, R. K., Hopkins, P. F., et al. 2025, The Astrophysical Journal, Volume 980, Issue 1, id.135, 15 pp., 980, 135
- Radcliffe et al. (2021) Radcliffe, J. F., Barthel, P. D., Garrett, M. A., et al. 2021, Astronomy and Astrophysics, 649, L9
@@
- Radcliffe et al. (2018) Radcliffe, J. F., Garrett, M. A., Muxlow, T. W., et al. 2018, Astronomy and Astrophysics, 619, A48
@@
- Rickard & Harvey (1984) Rickard, L. J. & Harvey, P. M. 1984, Astronomical Journal, Vol. 89, p. 1520-1530 (1984), 89, 1520
- Rivera et al. (2017) Rivera, G. C., Williams, W. L., Hardcastle, M. J., et al. 2017, Monthly Notices of the Royal Astronomical Society, 469, 3468
- Sarbadhicary et al. (2021) Sarbadhicary, S. K., Tremou, E., Stewart, A. J., et al. 2021, The Astrophysical Journal, 923, 31
- Sargent et al. (2010) Sargent, M. T., Schinnerer, E., Murphy, E., et al. 2010, The Astrophysical Journal Supplement Series, 186, 341
@@
- Schober et al. (2023) Schober, J., Sargent, M. T., Klessen, R. S., & Schleicher, D. R. 2023, Astronomy & Astrophysics, Volume 679, id.A47, ¡NUMPAGES¿20¡/NUMPAGES¿ pp., 679, A47
- Scoville et al. (2007) Scoville, N., Abraham, R. G., Aussel, H., et al. 2007, The Astrophysical Journal Supplement Series, 172, 38
@@
- Smith et al. (2021) Smith, D. J., Haskell, P., Gürkan, G., et al. 2021, Astronomy and Astrophysics, 648, A6
- Ulvestad et al. (2004) Ulvestad, J. S., Antonucci, R. R. J., & Barvainis, R. 2004, The Astrophysical Journal, Volume 621, Issue 1, pp. 123-129., 621, 123
@@
- van der Wel et al. (2021) van der Wel, A., Bezanson, R., D’Eugenio, F., et al. 2021, The Astrophysical Journal Supplement Series, 256, 44
- Wang et al. (2022) Wang, T. M., Magnelli, B., Schinnerer, E., et al. 2022, Astronomy and Astrophysics, 660, A142
- Whittam et al. (2017) Whittam, I. H., Jarvis, M. J., Green, D. A., Heywood, I., & Riley, J. M. 2017, Monthly Notices of the Royal Astronomical Society, 471, 908
@@
- Yoon (2024) Yoon, I. 2024, The Astrophysical Journal, 975, 15

### v2

@@ 1 Introduction
Several studies (Magnelli et al., 2015; Rivera et al., 2017; Delhaize et al., 2017) observed a dependence of the IRRC on the redshift, in the form of with . Other works (Gürkan et al., 2018; Smith et al., 2021; Delvecchio et al., 2021, D21 hereafter) have recently argued that the IRRC evolves primarily with , with more massive galaxies showing a systematically lower , while the dependence on the redshift is secondary and weaker, as indeed at fixed stellar mass (see also, Gürkan et al., 2018; Smith et al., 2021). Among the possible physical drivers of the decline observed in massive star-forming galaxies, the energy loss of cosmic ray electrons (CRe) has been suggested to play a significant role (e.g. Schober et al., 2023; Yoon, 2024). Specifically, Yoon (2024) modelled the CRe energy-loss as a function of the gas density and redshift of the host galaxy. Their findings indicate an anti-correlation between these quantities: specifically, galaxies with higher gas densities (or, equivalently, higher stellar masses, e.g. Kennicutt, 1998), experience less intense energy losses, arguably because they retain more easily their CRes. On the reverse, less massive galaxies are generally most affected by CRe’s energy losses. As a consequence of these synchrotron energy losses, less massive galaxies have lower radio luminous with respect to more massive ones.
@@
where is the observed density flux at the frequency and and are the major and minor axes of the beam (Ulvestad et al., 2005).
@@ 3 Observation and data calibration
where is the number of ’real’ (input) blobs and is the number of the (output) detected blobs. We count the output blobs within a radius of 0.2′′ from the input (real) source, similarly to what we do for real detections (see Sect.4). We find that the spuriousness is higher than 50% at S/N 5, while it decreases to 15% at S/N5.5, and exponentially drops below 2% at S/N6. If counting all random blobs over the full image (44′′), the spuriousness at fixed S/N increases proportionally to the searching area, except at S/N6 where it drops to below 2% as found in the central area.
4 Final sample of VLBA detections
@@
Then, we also find that the real distribution exhibits a deviation from the Gaussianity (e.g., when compared to the mirror distribution) on the left side, in both panels. However, this is minimal, as indeed a K-S test could not reject the hypothesis that these two distributions come from the same parent sample (with a p-value of 0.997).
@@ 7.1 Cumulative Distribution of AGN in the faintest regime
To do so, we convert the Euclidean-normalized 1.4 GHz source counts into a dimensionless quantity, re-scaled to the effective sky area of the (,rms)-matched CS (i.e. 9.7 x 10-6 sr). This yields a number of expected AGN in the same area, as a function of 1.4 GHz flux. We make sure that the relation between differential NAGN and S1.4 yields an integrated = 4 in the range 36–58 Jy (i.e. matching our VLBA detections, red star in Fig. 4). Then we extrapolate the same trend down, and below our nominal VLBA flux density limit. Finally, we compute the cumulative distribution of at Jy, and we found out that 42 sources would be cumulated at a flux density of 5 Jy, which is then the minimum flux density reached by the 42 brightest AGN. We stress again that this is an over-conservative assumption, since it is very unlikely that our non-detections are all clustered just below the flux limit. However, at these low flux densities, the brightness temperature drops to K (see Eq. 2). This suggests that, even if we were able to detect such faint sources with the VLBA, the origin of their emission would be ambiguous between AGN or SF. Consequently, we conclude that pushing VLBA observation to sub-Jy sensitivity would not be an effective strategy for revealing a residual hidden population of radio AGN. For instance, according to the ngVLA exposure time calculator 666https://ngect.nrao.edu/, we note that Band 1 (2.3 GHz observations) would reach brightness temperatures T K (at 5, i.e. 5 Jy/beam) in about 16h on-source. Hence, fainter fluxes, as expected for the bulk of our VLBA non-detections, would be detected with a lower T K, thus carrying an ambiguous origin despite their easy detection.
@@ 7.2 Diffuse AGN emission as a possible caveat
Assuming that this radio-excess traces the full AGN-related flux over the entire galaxy size, the comparison between total VLBA and AGN-driven VLA flux indicates how much flux is lost on larger scales. We note (see Figure 11 in Herrera-Ruiz et al., 2017) that at fluxes S1.4 , the VLBA flux is already of the total VLA AGN flux, hence highly compact. By extrapolating this trend towards fainter flux regimes (as for our VLBA targets), we would expect that the VLBA flux is capturing the vast majority (if not all) of the total AGN flux.
@@ Acknowledgements.
The authors thanks the referee for a constructive report that helped us clarify the results and implications of this work. G. P., I. D., M. G., F. U., and I. P. acknowledge funding by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006. C. S. acknowledges financial support from INAF Ricerca Fondamentale 2024 (Ob.Fu. 1.05.24.07.04), from the Italian Ministry of Foreign Affairs and International Cooperation (grant PGRZA23GR03), and by the Italian Ministry of University and Research (grant FIS-2023-01611, CUP C53C25000300001). IP also acknowledges support from INAF under the Large Grant 2022 funding scheme (project ”MeerKAT and LOFAR Team up: a Unique Radio Window on Galaxy/AGN co-Evolution”). W.W. acknowledges the grant support through JWST programs. Support for programs JWST-GO-03045 and JWST-GO-03950 were provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This research has also been supported by the European Regional Development Fund under grant agreement PK.1.1.10.0007 (DATACROSS). This work is based, in part, on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST.References
@@
- Algera et al. (2020) Algera, H. S. B., Smail, I., Dudzevičiūtė, U., et al. 2020, ApJ, 903, 138
- Appleton et al. (2004) Appleton, P. N., Fadda, D. T., Marleau, F. R., et al. 2004, ApJS, 154, 147
- Basu et al. (2015) Basu, A., Wadadekar, Y., Beelen, A., et al. 2015, ApJ, 803, 51
- Bell (2003) Bell, E. F. 2003, ApJ, 586, 794
- Bourne et al. (2011) Bourne, N., Dunne, L., Ivison, R. J., et al. 2011, MNRAS, Volume 410, Issue 2, pp. 1155-1173., 410, 1155
- Casey et al. (2023) Casey, C. M., Kartaltepe, J. S., Drakos, N. E., et al. 2023, ApJ, 954, 31
@@
- Condon (1992) Condon, J. J. 1992, ARA&A, Vol. 30, p. 575-611 (1992), 30, 575
@@
- Deane et al. (2024) Deane, R. P., Radcliffe, J. F., Njeri, A., et al. 2024, MNRAS, Volume 529, Issue 3, pp.2428-2442, 529, 2428
@@
- Delvecchio et al. (2022) Delvecchio, I., Daddi, E., Sargent, M. T., et al. 2022, A&A, 668, A81
- Delvecchio et al. (2021) Delvecchio, I., Daddi, E., Sargent, M. T., et al. 2021, A&A, 647, A123
- Delvecchio et al. (2017) Delvecchio, I., Smolčić, V., Zamorani, G., et al. 2017, A&A, 602, A3
@@
- Galluzzi et al. (2025) Galluzzi, V., Behiri, M., Giulietti, M., & Lapi, A. 2025, Galaxies, 13, 34
@@
- Gim et al. (2019) Gim, H. B., Yun, M. S., Owen, F. N., et al. 2019, ApJ, 875, 80
- Gürkan et al. (2018) Gürkan, G., Hardcastle, M. J., Smith, D. J., et al. 2018, MNRAS, 475, 3010
- Hales et al. (2012) Hales, C. A., Murphy, T., Curran, J. R., et al. 2012, MNRAS, 425, 979
- Harish et al. (2025) Harish, S., Kartaltepe, J. S., Liu, D., et al. 2025, ApJ, 992, 45
- Harwit & Pacini (1975) Harwit, M. & Pacini, F. 1975, ApJ, vol. 200, Sept. 15, 1975, pt. 2, p. L127-L129., 200, L127
- Helou & Rowan-Robinson (1985) Helou, G. & Rowan-Robinson, M. 1985, ApJ, 298, 7
- Herrera-Ruiz et al. (2017) Herrera-Ruiz, N., Middelberg, E., Deller, A., et al. 2017, A&A, 607, A132
- Herrera-Ruiz et al. (2018) Herrera-Ruiz, N., Middelberg, E., Deller, A., et al. 2018, A&A, 616, A128
- Heywood et al. (2022) Heywood, I., Rammala, I., Camilo, F., et al. 2022, ApJ, 925, 165
@@
- Ivison et al. (2010a) Ivison, R. J., Alexander, D. M., Biggs, A. D., et al. 2010a, MNRAS, 402, 245
- Ivison et al. (2010b) Ivison, R. J., Magnelli, B., Ibar, E., et al. 2010b, A&A, 518, L31
- Jarvis et al. (2010) Jarvis, M. J., Smith, D. J. B., Bonfield, D. G., et al. 2010, MNRAS, Volume 409, Issue 1, pp. 92-101., 409, 92
@@
- Jin et al. (2018) Jin, S., Daddi, E., Liu, D., et al. 2018, ApJ, 864, 56
- Kartaltepe et al. (2015) Kartaltepe, J. S., Mozena, M., Kocevski, D., et al. 2015, ApJS, 221, 11
- Kashino et al. (2019) Kashino, D., Silverman, J. D., Sanders, D., et al. 2019, ApJS, 241, 10
- Kennicutt (1998) Kennicutt, Jr., R. C. 1998, ARA&A, 36, 189
@@
- Madau & Dickinson (2014) Madau, P. & Dickinson, M. 2014, Annual Review of A&A, 52, 415
- Magnelli et al. (2015) Magnelli, B., Ivison, R. J., Lutz, D., et al. 2015, A&A, 573, A45
@@
- Middelberg et al. (2013) Middelberg, E., Deller, A. T., Norris, R. P., et al. 2013, A&A, 551, A97
@@
- Molnár et al. (2018) Molnár, D. C., Sargent, M. T., Delhaize, J., et al. 2018, MNRAS, 475, 827
- Molnár et al. (2021) Molnár, D. C., Sargent, M. T., Leslie, S., et al. 2021, MNRAS, 504, 118
@@
- Njeri et al. (2023) Njeri, A., Beswick, R. J., Radcliffe, J. F., et al. 2023, MNRAS, 519, 1732
@@
- Ponnada et al. (2025) Ponnada, S. B., Cochrane, R. K., Hopkins, P. F., et al. 2025, ApJ, Volume 980, Issue 1, id.135, 15 pp., 980, 135
- Radcliffe et al. (2021) Radcliffe, J. F., Barthel, P. D., Garrett, M. A., et al. 2021, A&A, 649, L9
@@
- Radcliffe et al. (2018) Radcliffe, J. F., Garrett, M. A., Muxlow, T. W., et al. 2018, A&A, 619, A48
@@
- Rickard & Harvey (1984) Rickard, L. J. & Harvey, P. M. 1984, AJ, Vol. 89, p. 1520-1530 (1984), 89, 1520
- Rivera et al. (2017) Rivera, G. C., Williams, W. L., Hardcastle, M. J., et al. 2017, MNRAS, 469, 3468
- Sarbadhicary et al. (2021) Sarbadhicary, S. K., Tremou, E., Stewart, A. J., et al. 2021, ApJ, 923, 31
- Sargent et al. (2010) Sargent, M. T., Schinnerer, E., Murphy, E., et al. 2010, ApJS, 186, 341
@@
- Schober et al. (2023) Schober, J., Sargent, M. T., Klessen, R. S., & Schleicher, D. R. 2023, Astronomy & Astrophysics, Volume 679, id.A47, 20 pp., 679, A47
- Scoville et al. (2007) Scoville, N., Abraham, R. G., Aussel, H., et al. 2007, ApJS, 172, 38
@@
- Smith et al. (2021) Smith, D. J., Haskell, P., Gürkan, G., et al. 2021, A&A, 648, A6
- Ulvestad et al. (2005) Ulvestad, J. S., Antonucci, R. R. J., & Barvainis, R. 2005, ApJ, Volume 621, Issue 1, pp. 123-129., 621, 123
@@
- van der Wel et al. (2021) van der Wel, A., Bezanson, R., D’Eugenio, F., et al. 2021, ApJS, 256, 44
- Wang et al. (2022) Wang, T. M., Magnelli, B., Schinnerer, E., et al. 2022, A&A, 660, A142
- Whittam et al. (2017) Whittam, I. H., Jarvis, M. J., Green, D. A., Heywood, I., & Riley, J. M. 2017, MNRAS, 471, 908
@@
- Yoon (2024) Yoon, I. 2024, ApJ, 975, 15